# MVP-Shapley: Feature-based Modeling for Evaluating the Most Valuable Player in Basketball

### v2

@@ Abstract
Index Terms:
Shapley Value, MVP Evaluation, Basketball
@@ II Preliminaries
FGField Goal Percentage
@@
3P3-Point Field Goal Percentage
@@
FTFree Throw Percentage
@@
TSTrue Shooting Percentage
eFGEffective Field Goal Percentage
@@
ORBOffensive Rebound Percentage
DRBDefensive Rebound Percentage
TRBTotal Rebound Percentage
ASTAssist Percentage
STLSteal Percentage
BLKBlock Percentage
TOVTurnover Percentage
USGUsage Percentage
@@ III-A Feature Preprocessing
Assuming the maximum number of players on each team is $p$ with any missing players padded with zeros, each player’s statistical data is represented by a vector of dimension $q$ (including basic and advanced statistics). For each match, the features of all players from both teams are concatenated into a feature vector of dimension $2pq$. The first $pq$ dimensions correspond to the home team’s players, followed by the $pq$ dimensions representing the away team’s players. The label denotes the home team’s game result, with 1 representing a home team win and 0 representing a home team loss. It is crucial to understand that "home team" refers to the team whose features are presented first in the data. For each pair of teams in a game, the outcomes for the winning and losing teams are complementary, enabling the construction of two data points with opposite labels. For instance, if team A defeats team B in a game, let the sets of all players from team A and team B be $\{a_{1},\cdots,a_{p}\}$ and $\{b_{1},\cdots,b_{p}\}$, respectively. The statistical data features of player $a_{i}$ are recorded as $\{a_{i1},\cdots,a_{iq}\}$, and those of player $b_{i}$ as $\{b_{i1},\cdots,b_{iq}\}$. Consequently, two data points are constructed for this game:$\{(x_{1},y_{1}),(x_{2},y_{2})\}$:
@@ III-E Refinements in Algorithmic Approach
[…]
@@ IV-A Experimental Setup
During the training of the win-loss model, we split the dataset into a training set and a test set with a 9:1 ratio. The models trained on the NBA and Dunk City Dynasty datasets achieved accuracies of 0.99 and 0.97, respectively. Here, we employ TreeSHAP [23] to compute the Shapley values efficiently. For the NBA dataset, we divide the variables into five groups according to their importance: ‘+/-’, ‘ORtg’, ‘DRtg’, ‘BPM’, and the remaining variables. MVP-Shapley has three versions with different utility functions. Different versions train the win-loss models by constructing different features. The first version (Ours_$\mathbf{V_{1}}$) uses all features. The second version (Ours_$\mathbf{V_{2}}$) uses the remaining features after removing ‘+,-’ and ‘DRtg’. The third version (Ours_$\mathbf{V_{3}}$) uses the features after fuzzification. For ‘+/-’, we set $|B|=3$ and for ‘DRtg’, $|B|=8$. Here, the second version focuses solely on combinatorial optimization, while the third version incorporates fuzzy optimization. For the Dunk City Dynasty dataset, we use the optimized features for analysis in Section IV-C3.
@@ IV-A2 Evaluation Metrics
1) Average Rank Difference (ARD): The average rank difference quantifies the average discrepancy between the predicted MVP rank and the actual rank of the corresponding player. Mathematically, ARD is calculated as follows: $ARD=\frac{1}{N}\sum_{i=1}^{N}|rank_{\text{predicted MVP}_{i}}-rank_{\text{ground truth}_{i}}|.$ $N$ is the total number of players on the MVP voting list. $rank_{\text{predicted MVP}_{i}}$ is the rank of the predicted MVP for the i-th player on the MVP voting list, $rank_{\text{ground truth}_{i}}$ is the actual voting rank of the i-th player.
@@ IV-C1 NBA Regular Season MVP Results
We analyzed the outcomes of the regular season MVP experiments over the past three years. $\mathbf{M_{1}}$, $\mathbf{M_{2}}$, and $\mathbf{M_{3}}$ correspond to the MVP evaluation method in Section III-D2. We tested three evaluation metrics of these methods on three different versions of the win-loss model. TABLE V illustrates the comparative results of different methods for all the players ranked in the MVP voting. Table VI presents the comparative results of different methods for all the players ranked in the top three in the MVP voting. We can draw the following conclusions: First, our methods outperform the baselines. Second, the second version (Ours_$V_{2}$) of our methods is particularly effective in evaluating the regular season MVP. TABLE IV shows the MVP selection results of various methods for the 2023/2024 season. $ACRW$ represents the average player’s contribution ranking across all winning games. $ACRA$ represents the average player’s contribution ranking across all games. $AC$ represents the average player’s contribution across all games. Our methods all successfully selected the MVP of the season except $M_{1}$ (Ours_$V_{3}$).
@@
$M_{1}$ (Ours_$V_{1}$) 9.00 0.93 0.56 25.00 0.41 0.31 23.42 0.39 0.33
$M_{2}$ (Ours_$V_{1}$) 15.72 0.82 0.33 44.42 0.54 0.31 39.17 0.62 0.58
$M_{3}$ (Ours_$V_{1}$) 14.00 0.72 0.33 44.54 0.58 0.31 34.50 0.35 0.58
$M_{1}$ (Ours_$V_{2}$) 17.89 0.68 0.67 11.85 0.57 0.54 19.92 0.32 0.50
$M_{2}$ (Ours_$V_{2}$) 12.56 0.76 0.56 8.54 0.67 0.62 23.08 0.42 0.50
$M_{3}$ (Ours_$V_{2}$) 6.22 0.85 0.67 5.92 0.59 0.77 9.25 0.43 0.67
$M_{1}$ (Ours_$V_{3}$) 26.22 0.58 0.22 33.92 0.68 0.46 39.75 0.56 0.33
$M_{2}$ (Ours_$V_{3}$) 23.89 0.72 0.33 47.77 0.78 0.23 44.17 0.58 0.42
$M_{3}$ (Ours_$V_{3}$) 20.67 0.77 0.33 48.54 0.68 0.31 36.17 0.74 0.42
@@
$M_{1}$ (Ours_$V_{1}$) 0.33 1.00 0.67 9.67 0.50 0.33 6.67 -0.50 0.33
$M_{2}$ (Ours_$V_{1}$) 3.00 1.00 0.67 1.67 0.50 0.67 4.00 0.50 0.33
$M_{3}$ (Ours_$V_{1}$) 4.67 1.00 0.67 0.67 0.50 1.00 6.67 -0.50 0.00
$M_{1}$ (Ours_$V_{2}$) 2.00 1.00 0.67 2.67 0.50 0.67 1.33 -0.50 1.00
$M_{2}$ (Ours_$V_{2}$) 0.67 1.00 0.67 2.33 0.50 0.67 0.00 1.00 1.00
$M_{3}$ (Ours_$V_{2}$) 0.67 0.50 0.67 2.0 0.50 0.67 0.67 0.50 1.00
$M_{1}$ (Ours_$V_{3}$) 10.33 1.00 0.33 4.67 -0.50 0.33 2.00 -0.50 0.67
$M_{2}$ (Ours_$V_{3}$) 5.33 1.00 0.33 0.67 0.50 1.00 1.00 -0.50 0.67
$M_{3}$ (Ours_$V_{3}$) 4.00 1.00 0.33 1.67 -1.00 0.67 1.00 1.00 0.67
@@ IV-C2 NBA Finals MVP Results
$M_{1}$ (Ours_$V_{1}$) 6 2 2 4 5 3 1 4 1 2
$M_{2}$ (Ours_$V_{1}$) 6 3 3 1 4 4 1 4 3 1
$M_{3}$ (Ours_$V_{1}$) 1 3 3 2 2 4 1 3 3 1
$M_{1}$ (Ours_$V_{2}$) 5 1 3 1 1 1 1 1 1 1
$M_{2}$ (Ours_$V_{2}$) 5 1 1 1 1 1 1 1 1 1
$M_{3}$ (Ours_$V_{3}$) 5 1 1 1 1 1 1 1 1 1
$M_{1}$ (Ours_$V_{3}$) 2 1 2 1 1 2 1 2 1 1
$M_{2}$ (Ours_$V_{3}$) 2 1 1 1 1 1 1 2 1 1
$M_{3}$ (Ours_$V_{3}$) 2 1 1 1 1 1 1 1 1 3
@@ IV-C5 Online Service Deployment Framework for MVP-Shapley
For the performance comparison with the original method, we conducted an online A/B test for over one month to compare player report rates ($RR=\frac{N_{\text{reports}}}{N_{\text{players}}}\times 100\%$) and churn rates ($CR=\frac{N_{\text{churned}}}{N_{\text{start\_players}}}\times 100\%$). Where $N_{\text{reports}}$ = Number of reports received, $N_{\text{players}}$ = Total number of active players during the same period, $N_{\text{churned}}$ = Number of players who stopped playing the game, $N_{\text{start\_players}}$ = Total number of active players at the beginning of the period. We employed a rigorous hash-based random bucketing method for our A/B test, with 10,000 players assigned to each bucket. Our method has reduced the report rate by approximately $9.64\%\pm 1.2\%$ and the churn rate by approximately $8.11\%\pm 0.9\%$ compared to the existing method. Experimental results indicate the superiority of our method.
@@ IV-C6 Computational Efficiency of Shapley Values
  • $<$5 minutes for full season NBA analysis
IV-D Feature Visualization
Fig. 6 and Fig. 7 respectively demonstrate the change in feature importance before and after optimization for the NBA dataset and the Dunk City Dynasty dataset. Fig. 8 illustrates the impact of different features on the outcome of actual games for Dunk City Dynasty.

### v3

@@ II Preliminaries
FG% Field Goal Percentage
@@
3P% 3-Point Field Goal Percentage
@@
FT% Free Throw Percentage
@@
TS% True Shooting Percentage
eFG% Effective Field Goal Percentage
@@
ORB% Offensive Rebound Percentage
DRB% Defensive Rebound Percentage
TRB% Total Rebound Percentage
AST% Assist Percentage
STL% Steal Percentage
BLK% Block Percentage
TOV% Turnover Percentage
USG% Usage Percentage
@@ III-A Feature Preprocessing
[…]
@@ III-E Refinements in Algorithmic Approach
In our analysis, we possess the ground truth of the MVP voting data with various statistical metrics related to players. We process the statistical data into features to train a win-loss model and compute the utility function. However, we are uncertain about which features are the true causal factors leading to the ground truth results. Therefore, our approach involves seeking superior causal factors from the outcomes. Let $Y$ represent the ground truth, and $X_{1},X_{2},\ldots,X_{n}$ denote the variables related to computing the utility function before feature preprocessing. Our objective is to determine the optimal set of variables $X^{*}$ such that: $X^{*}=\underset{X}{\arg\max}\,\mathbb{E}[Y|X]$, where $\mathbb{E}[Y|X]$ denotes the conditional expectation of the ground truth given the set $X$. However, the computational complexity is excessively high with $2^{n}$ possible combinations. To address this, we propose using feature importance ranking to group features with low importance together, thereby reducing the complexity of variable combinations. For example, Fig. 6 illustrates the influence of different features on win-loss model predictions of the NBA dataset. The top two most influential variables, ‘+,-’ and ‘DRtg’ can be considered as individual variables, and all the remaining variables can be considered as a whole. Then we have only 8 possible combinations.
@@ IV-A Experimental Setup
[…]
@@ IV-A2 Evaluation Metrics
1) Average Rank Difference (ARD): The average rank difference quantifies the average discrepancy between the predicted MVP rank and the actual rank of the corresponding player. Mathematically, ARD is calculated as follows: $ARD=\frac{1}{N}\sum_{i=1}^{N}|rank_{\text{predicted MVP}_{i}}-rank_{\text{ground truth}_{i}}|.$ NN is the total number of players on the MVP voting list. $rank_{\text{predicted MVP}_{i}}$ is the rank of the predicted MVP for the i-th player on the MVP voting list, $rank_{\text{ground truth}_{i}}$ is the actual voting rank of the i-th player.
@@ IV-C1 NBA Regular Season MVP Results
[…]
@@
[…]
@@
[…]
@@ IV-C2 NBA Finals MVP Results
[…]
@@ IV-C5 Online Service Deployment Framework for MVP-Shapley
For the performance comparison with the original method, we conducted an online A/B test for over one month to compare player report rates ($RR=\frac{N_{\text{reports}}}{N_{\text{players}}}\times 100\%$) and churn rates ($CR=\frac{N_{\text{churned}}}{N_{\text{start\_players}}}\times 100\%$). Where $N_{\text{reports}}$ = Number of reports received, $N_{\text{players}}$ = Total number of active players during the same period, $N_{\text{churned}}$ = Number of players who stopped playing the game, $N_{\text{start\_players}}$ = Total number of active players at the beginning of the period. We employed a rigorous hash-based random bucketing method for our A/B test, with 10,000 players assigned to each bucket. Our method has reduced the report rate by approximately $9.64\text{\textperthousand}\pm 1.2\text{\textperthousand}$ and the churn rate by approximately $8.11\text{\textperthousand}\pm 0.9\text{\textperthousand}$ compared to the existing method. Experimental results indicate the superiority of our method.
@@ IV-C6 Computational Efficiency of Shapley Values
  • <<5 minutes for full season NBA analysis
IV-D Feature Visualization Analysis
Fig. 6 and Fig. 7 respectively demonstrate the change in feature importance before and after optimization for the NBA dataset and the Dunk City Dynasty dataset. Fig. 6 shows the feature importance without any processing, where the ‘+/-’ feature dominates the results, making it difficult for other features to contribute to the interpretation—an unreasonable scenario. After removing such features and applying fuzzification, Fig. 6 reveals a more balanced influence of different features on the results, enabling each feature to contribute meaningfully to the interpretation. Fig. 8 illustrates the impact of different features on the outcome of actual games for Dunk City Dynasty. These results can help players analyze their performance and serve as a personalized coach to improve specific skills.